OpenStudy (anonymous):

Simplify the following expression square root sign with -104 in it and square root sign with only a - sign in it. please please can someone help me with this one?

OpenStudy (anonymous):

U can use sqrt(ab) = sqrt(a)*sqrt(b)to simplify sqrt(104). The other seems to be some kind of mistake.

OpenStudy (anonymous):

oh i am sorry it is -8

OpenStudy (anonymous):

in the second square root

OpenStudy (anonymous):

can you work this one step by step for me please so i can see how to do

OpenStudy (anonymous):

Oh, they are both minus so take out a square root of -1 first of all and then use sqrt(ab) = sqrt(a)*sqrt(b)

OpenStudy (anonymous):

show me how to do this please. this is very new to me

OpenStudy (anonymous):

thank you for your help

OpenStudy (anonymous):

Do you know what "i" is?

OpenStudy (anonymous):

refresh my memory please

OpenStudy (anonymous):

i is the square root of -1 If you don't know that you probably shouldn't be attempting this kind of question.

OpenStudy (anonymous):

i have to it is one of my problems

OpenStudy (anonymous):

i know what square root means, however what is the square root of -1

OpenStudy (anonymous):

i

OpenStudy (anonymous):

if you could just help me solve the problem i posted

OpenStudy (anonymous):

I thought that I was doing that but you appear to have no idea what I am talking about.

OpenStudy (anonymous):

Perhaps someone else will help u.

OpenStudy (anonymous):

ok so sorry to bother you

OpenStudy (cwrw238):

square root of -1 is written as i so sqrt(-104) = sqrt 104 * i sqrt 104 = sqrt(4*26) = sqrt4 * sqrt26 = 2 sqrt26 sqrt(-104) = 2i sqrt26

Directrix (directrix):

@connielynn Is this a multiplication or addition problem?